# Anisotropic Spin Ice on a Breathing Pyrochlore Lattice

### v1

@@ Appendix A Experimental considerations
We now discuss possible realizations of the presented toy model. The system Hamiltonian (3) is based on the classical spin ice model. Due to large crystal field anisotropies, the spins are constrained to point along local axes connecting each corner point of the tetrahedron to the midpoint of the latter. This simplifies the Hamiltonian to an Ising-type interaction Hamiltonian between all spins on a single tetrahedron [2, 1]. The energy of this system is minimized when the spin-sum on all tetrahedra on this lattice vanishes. However, this simple ground-state behavior in spin ice can change under the influence of strain, pressure, or when an external field is applied [35, 69, 36, 17, 19]. Pressure along certain crystal axes may reduce the distance between the crystal sites along this direction, causing the ionic wave function to overlap and the exchange interaction between the neighbors to change [69], such that $H_{\mathrm{strain}}=\sum_{\braket{ij}}J_{ij}S_{i}^{z}S_{j}^{z}$. In an experimental study on $\mathrm{Dy}_{2}\mathrm{Ti}_{2}\mathrm{O}_{7}$, it was shown that pressure along the $[110]$ direction can lead to a splitting of the six ground states into a set of four lower lying and two higher lying states [35].
In a potential breathing pyrochlore system, we could also assume pressure to be applied along certain crystal axes and additionally suppose that the compression of the two tetrahedral sublattices differ from each other. The compression generally depends on the interaction strength between the sites within a crystal [70]. In the breathing pyrochlore, it is known that the interaction strength differs between sublattices which motivates our assumption of different compression on the individual sublattices. In our investigated toy model, we assume the interaction between the spin pairs $\{S_{0}^{z},S_{3}^{z}\}$ and $\{S_{1}^{z},S_{2}^{z}\}$ to differ when pressure is applied along $[001]$ or on the plane perpendicular to this direction.
Moreover, it was shown in different spin-ice thin films that epitaxial strain on the lattice can be generated by growing spin-ice compounds on a substrate with a different lattice constant. Depending on the growth direction of the spin ice compound on the growing substrate, the thin film can be strained in different directions [19, 20, 30]. A breathing pyrochlore structure of the substrate can thus lead to distinct strain on the individual tetrahedral sublattices, leading to the desired spin couplings.
To summarize, the requirements for a material to fulfill this behavior are the following: a (breathing) pyrochlore lattice is needed where the interactions in the lattice—on both sublattices—match the Hamiltonian given in (3). To get the anisotropic Hamiltonian, we need, for example, a lattice with different compressions for the individual tetrahedral sublattices or a growing substrate with a difference in bond length between the sublattices.
Appendix B Local spin basis
@@
$\begin{split}\mathbf{R}_{0}&=(100),\qquad\mathbf{R}_{1}=(1\frac{1}{2}\frac{1}{2}),\\ \mathbf{R}_{2}&=(\frac{1}{2}\frac{1}{2}0),\qquad\mathbf{R}_{3}=(\frac{1}{2}0\frac{1}{2}).\\ \end{split}$ (10)
@@
$\begin{split}\mathbf{r}_{0}&=\frac{1}{4}(100),\qquad\mathbf{r}_{1}=\frac{1}{4}(001),\\ \mathbf{r}_{2}&=\frac{1}{4}(111),\qquad\mathbf{r}_{3}=\frac{1}{4}(010).\\ \end{split}$ (11)
@@
Appendix C Self-consistent Gaussian approximation
@@
To calculate this and the structure factor in the large-$N$ limit, we need to perform the Fourier transform of the Hamiltonian (13). This gives
@@
|$\braket{}{\sigma_{n}(\mathbf{q})}{{}^{2}}=\frac{1}{\lambda+\beta\varepsilon_{n}(\mathbf{q})}.$ (22)
@@
We can now discuss some limiting cases for $\lambda$: At high temperatures, $\beta\rightarrow 0$, all energy modes are equally singly occupied and $\lambda=1$. For the isotropic nearest neighbor spin ice, as $\beta\rightarrow\infty$, the lowest two energy modes become flat and fully degenerate and $\lambda\rightarrow\frac{1}{2}$. This tells us that the high energy modes depopulate and the degenerate low energy modes are doubly occupied. For the planar ice, as $\beta\rightarrow\infty$, only one lowest energy mode becomes flat and $\lambda\rightarrow\frac{1}{4}$. If a normal mode becomes critical at any temperature, then $\lambda^{-1}\rightarrow\infty$ or $\lambda\rightarrow 0$. This indicates a phase transition with long-range order in the large-$N$ approximation.
Figure 6: Lagrange-multiplier for SCGA. Lagrange-multiplier that fixed the spin constraint in the self-consistent Gaussian approximation. For isotropic spin ice, $\lambda\rightarrow\frac{1}{2}$ as $T\rightarrow 0$. For planar ice, $\lambda\rightarrow\frac{1}{4}$ as $T\rightarrow 0$. For the line-order phase, $\lambda\rightarrow 0$ as $T\rightarrow 0$, suggesting long-range order at $T=0$. For the three remaining phases, the sudden drop of $\lambda\rightarrow 0$ at a finite temperature indicates a finite-temperature phase transition.
The Lagrange multiplier for the six phases is shown in Fig. 6. We can observe the expected behaviour as $T\rightarrow 0$ for isotropic spin ice ($\lambda\rightarrow\frac{1}{2}$) and for planar ice ($\lambda\rightarrow\frac{1}{4}$). For the line-order phase, we can observe how $\lambda$ monotonically decreases as $T$ approaches 0. This suggests a transition to long-range order to appear only at $T=0$, similarly to what is expected in a one-dimensional Ising-chain [71]. By contrast, the Lagrange-multiplier of omni-plane phase, the plane-paramagnetic phase, and the ferromagnetic phase seem to have a sudden drop to zero at a finite temperature. It is important to note, however, that $\lambda$ can only really reach zero at a finite temperature for $L\rightarrow\infty$. While for the plane-paramegnetic phase and the ferromagnetic phase, the drop in $\lambda$ appears sudden and steep, $\lambda$ of the omni-planar phase approaches small values more gradually. A finite-size scaling analysis of $\lambda$ in the omni-planar phase, however, still points towards a finite-temperature phase transition within the large-$N$ approximation. This supports the numerical findings in Sec. IV.2, indicating a finite-temperature phase transition in the three phases with any $\delta_{i}<0$.
@@
Appendix D Monte-Carlo simulation
@@
During each Monte-Carlo time step, six types of updates were employed to enhance equilibration: single spin flips, worm loop updates, and other cluster-updates (see App. E). For each measurement temperature, the system was evolved until $N_{\text{sweeps}}=N_{\text{spin}}\cdot 10^{3}$ accepted updates (including spin flips and worm loops) were performed. Measurements were taken at equilibrium.
To generate data across different temperatures, the system was progressively cooled in 10 steps from the last measurement temperature. At each temperature step, the system was equilibrated before proceeding to the next step. Finally, at the next measurement-temperature, an additional equilibration phase was conducted. Here again, the next sweep measurement was performed.
Appendix E Cluster update schemes
For isotropic nearest-jneighbor spin ice, in numerical simulations, typically updates are given by single spin flips as well as worm loop updates [41, 42, 43]. Since at low temperatures, spin freezing makes equilibration times with only single-spin flips extremely long, the worm loop algorithm finds update moves at a zero energy cost [14]. The ground state degeneracy of all anisotropic systems is, however, greatly reduced compared to the isotropic spin ice case. This also leads to an increase in equilibration time in numerical simulations, when only employing the two aforementioned update schemes. The problem lies in the acceptance rate of the attempted updates. A measurement series at a particular temperature is only completed if a prefixed number of spin updates are accepted. For isotropic spin ice below the crossover, where the system is in a ground state configuration, all attempted worm loop updates are accepted. However, this acceptance rate drops significantly for all anisotropic systems. Now, to achieve the same number of accepted updates, either the number of attempted updates must be increased significantly, or a new type of update with a higher acceptance rate must be proposed. The latter is oped for to produce the data in Fig. 3 and the particular update schemes are discussed here.
@@ Appendix E Cluster update schemes
The underlying ordering processes are not changed by the introduction of the new cluster-update schemes, but rather, achieving the result numerically is merely accelerated. This is shown exemplarily in Fig. 8 for a system size of $L_{x}\times L_{y}\times L_{z}=8$, where the data obtained by only using the worm loop updates and single spin flips, and all six cluster update schemes detailed above are shown. The two curves overlapp fully.

### v2

@@ Appendix A Experimental considerations
In this minimal model, the ground-state condition is that the net spin sum on each tetrahedron vanishes. However, this simple ground-state behavior in spin ice can change under the influence of strain, pressure, or when an external field is applied [42, 82, 43, 17, 19]. Pressure along certain crystal axes may reduce the distance between the crystal sites along this direction, causing the ionic wave function to overlap and the exchange interaction between the neighbors to change [82], such that $H_{\mathrm{strain}}=\sum_{\braket{ij}}J_{ij}S_{i}^{z}S_{j}^{z}$. While the precise relationship between strain and the individual contributions $J_{\text{nn}}$ and $D_{\text{nn}}$ is material dependent and beyond the scope of our minimal model, it is known that dipolar interactions ($D\sim r^{-3}$) are significantly less sensitive to lattice distortions than exchange interactions ($J\sim r^{-q}$, with $q\gg 3$[83, 84]. Therefore, strain primarily modulates the exchange term.
Experimental support for this assumption comes from studies on single crystals of Dy2Ti2O7, where uniaxial pressure along the $[110]$ direction was shown to lift the six-fold degeneracy of spin ice, splitting it into a lower-energy set of four and a higher-energy set of two configurations [42]. This observation is consistent with the phenomenology captured by our model.
While in bulk spin ice one typically expects small perturbations, recent experiments on thin-film systems suggest that even sizeable modifications of the nearest-neighbor coupling might indeed be achievable. Growing spin-ice compound thin films on a substrate with a different lattice constant can hereby lead to epitaxial strain of the lattice. Depending on the growth direction of the spin ice compound on the growing substrate, the thin film can be strained in different directions [19, 20, 30]. In epitaxially strained thin films of Dy2Ti2O7, the strain effects have been estimated to shift $J_{\text{nn}}$ by 0.050.05K - 0.50.5[19], and the degeneracy of the six low-energy spin ice configurations has been estimated to split via the strain up to 0.550.55[20], corresponding to $4\delta$ in our model. This suggests the parameter range chosen in our simulations of Sec. IV ($\delta=0.05J$) is within an experimentally realistic range.
While these results were obtained in uniform pyrochlore systems, our work also extended to the breathing pyrochlore lattice, where bond strengths can differ between the two tetrahedral sublattices. In a potential breathing pyrochlore material, it is conceivable that pressure applied along specific crystal directions could compress the two types of tetrahedra differently. The compression generally depends on the interaction strength between the sites within a crystal [85]. In the breathing pyrochlore, it is known that the interaction strength differs between sublattices, which motivates our assumption of different compression on the individual sublattices. In our investigated toy model, we assume the interaction between the spin pairs $\{S_{0}^{z},S_{3}^{z}\}$ and $\{S_{1}^{z},S_{2}^{z}\}$ to differ when pressure is applied along $[001]$ or on the plane perpendicular to this direction. Another possible route to achieve the staggered strain on the different tetrahedral sublattices could involve using epitaxially strained thin films grown on a breathing pyrochlore substrate. If the lattice mismatch between film and substrate differs between the two types of tetrahedra, this might, in principle, induce strain of opposite sign on the two sublattices. While this approach would require significant material engineering, it offers a potential path toward realizing the asymmetry in $\delta$ considered in our model Hamiltonian.
To summarize, realizing the anisotropic breathing spin ice model experimentally would require a (breathing) pyrochlore lattice in which the interactions on both sublattices match those of the Hamiltonian (3). Achieving the desired anisotropic bond dependence may be experimentally feasible through several routes, including sublattice-selective compression of single crystal breathiong pyrochlores via application of directional pressure, or epitaxial strain engineering using lattice-mismatched substrates.
Appendix B Omni-Plane projections and transition temperatures
When $\delta_{A}<0$ and $\delta_{B}=0$, then a ground state for the system is given by all configurations that have $\mathsf{Q}_{A}=0$ and the partial charge $\mathsf{q}_{A}^{03}=-\mathsf{q}_{A}^{12}=\pm 2$, while on the $B$ sublattice only the total charge must be $\mathsf{Q}_{B}=0$; the value of the partial charge is not fixed. We find that this constraints leads to three types of ground state constructions: (a) in a plane perpendicular to the $z$ direction the $A$ tetrahedra are antiferromagnetically aligned, (b) in a plane perpendicular to the $y$ direction the $A$ tetrahedra are ferromagnetically aligned, or, (c) in a plane perpendicular to the $x$ direction the $A$ tetrahedra are ferromagnetically aligned. This structure is shown in Fig. 6(c-e) as projected onto the respective plane. Fulfilling requirement (a) means that in $z$ directions the alignment of planes is independent, but within $x,y$-planes, the tetrahedra are correlated. The counting of the total number of states is straightforward: we start from a configuration (a), (b), or (c) and count in how many different ways we can flip planes perpendicular to the specified direction. In direction $\alpha\in\{x,y,z\}$ there are $2L_{\alpha}$ individual planes of $A$ tetrahedra. Therfore there are $\binom{2L_{\alpha}}{i}$ possible ways to flip $i$ planes of $A$ tetrahedra. Counting all possibilities would give us the total number of configurations $\Omega=\sum_{\alpha=x,y,z}\sum_{i=1}^{2L_{\alpha}}\binom{2L_{\alpha}}{i}=\sum_{\alpha=x,y,z}2^{2L_{\alpha}}$. However, here, six states are counted twice as explained below:
There are only two states that at the same time fulfill requirement (a) and requirement (b): when in a plane perpendicular to the $z$ direction, the $A$ tetrahedra are antiferromagnetically aligned and the magnetization of neighboring planes is alternating. One of these states is shown in Fig. 6(f) with the respective projection onto the plane. There are only two states that at the same time fulfill requirement (a) and requirement (c): when in a plane perpendicular to the $x$ direction, the $A$ tetrahedra are ferromagnetically aligned and the magnetization of neighboring planes is alternating. One of these states is shown in Fig. 6(g) with the respective projection onto the plane. There are only two states that at the same time fulfill requirement (b) and requirement (c): when in a plane perpendicular to the $y$ direction, the $A$ tetrahedra are ferromagnetically aligned and all the planes’ magnetizations are equally aligned. One of these states is shown in Fig. 6(h) with the respective projection onto the plane. This, in total, gives six states that are counted twice when considering all possible configurations that can be generated by the plane flips. To account for this double counting, we subtract those states once from our calculations to reach $\Omega=\sum_{\alpha=x,y,z}\sum_{i=1}^{2L_{\alpha}-1}\binom{2L_{\alpha}}{i}=\sum_{\alpha=x,y,z}(2^{2L_{\alpha}}-2)$.
To understand the spin structure factor (SSF) shown in Fig. 4, we can imagine starting from any of the six overlapping ground states. The projections onto the three planes ($x,y$, $x,z$, and $y,z$) of the overlapping ground states are shown in Fig. 6(f-h). In the $x,y$-plane, we either have a horizontally or a vertically striped pattern on a square lattice. This would lead to Bragg peaks at $(0,\pm 1)$ or $(\pm 1,0)$ in the square lattice. In the $x,z$-plane and the $y,z$-plane, we either have ferromagnetic or antiferromagnetic alignment on a square lattice Brillouin zone. This would lead to a Bragg peak at $(0,0)$ or $(\pm 1,\pm 1)$ in the square lattice Brillouin zone. Flipping any plane of a tetrahedron starting from any of these states will introduce a line of defects into the square lattice projection in the plane perpendicular to the flipped plane. This is shown with the example of Fig.6(i). The line defects lead to a broadening of the Bragg peaks along the line-defect direction. This leads to an elongated peak at $(\pm 1,0,0)$ and $(0,\pm 1,0)$ from the stripe order in the $[hl0]$ (or $x,y$-) plane. On the other hand, the Bragg peak from the ferromagnetic order at $(0,0,0)$ is elongated horizontally, and the Bragg peaks from the antiferromagnetic order at $(\pm 1,0,\pm 1)$ are elongated vertically. This created the appearance of a square grid.
To estimate the critical temperature of the omni-plane and plane-paramagnetic phase, we employed a Peierls-type argument [86] based on the energy cost and entropy of defect loops.
We can first analyze the argument in the plane-paramagnetic phase. When the planes decouple, the $B$ tetrahedra form an antiferromagnetic square lattice. By forming octagon-shaped loops between four neighboring $B$ tetrahedra in a single plane, we can create four partial changes in the $B$ tetrahedra touched by the octagon, each with an energy difference of $\Delta E=4\delta$. An octagon loop connecting four tetrahedra in a common plane is shown in Fig.6(j). This is the equivalent of thinking about flipping a single spin (in the middle of the octagon) in the square lattice and creating four frustrated bonds (the four defects on the $B$ tetrahedra). Linking together octagons creates $L$ defects at the outermost boundary of the updated area with perimeter $L$. The possibilities to construct an updated area with perimeter $L$ may be named $\mu$. From this, we get
[…]
Figure 6: Omni-Plane ground state projections onto different planes. (a) The projection of a single conventional unit cell of the pyrochlore lattice onto the x,y-plane shows the four layers of tetrahedra as projected onto the plane. (b) The two ground-state configurations of an $A$ tetrahedron have a global magnetization vector pointing parallel or antiparallel to the global $z$ direction. This is shown by the projected tetrahedron-square having color blue or red. (c-e) Projection of the unit cell in a ground-state configuration onto different planes. Each layer of $A$ tetrahedra forms a square lattice, tilted by $45^{\circ}$ with respect to the unit cell. For the different ground-states in different planes, the resulting square lattice has AA tetrahedra being antiferromagnetically ($x,y$-plane) or ferromagnetically ($x,z$- and $y,z$-plane) aligned. (f-g) The three states that simultaneously fulfill two of the ground-state criteria in two perpendicular planes. The projection of the unit cell forms a square lattice with a stripe-pattern, an alternating-pattern, or a uniform pattern. On the right-hand side, the position of two AA-tetrahedra in different unit-cell projections is shown. (i) Flipping a plane of $A$ tetrahedra in the $y,z$-plane generates another ground state, but creates a line defect in the uniform pattern of the $x,y$-plane. (j) Flipping flips in an octagon-loop shape within a single tetrahedra plane, creates partial defects at the sides of the octagon. Grouping together octagons creates partial defects at the outermost border. (k) In the omni-plane phase the flipping-loops can also connect different planes.
Appendix C Local spin basis
@@
[…]
@@
[…]
@@
Appendix D Self-consistent Gaussian approximation
@@
To calculate this and the structure factor in the SCGA, we need to perform the Fourier transform of the Hamiltonian (13). This gives
@@
$\braket{|\sigma_{n}(\mathbf{q})|{}^{2}}=\frac{1}{\lambda+\beta\varepsilon_{n}(\mathbf{q})}.$ (22)
@@
We can now discuss some limiting cases for $\lambda$: At high temperatures, $\beta\rightarrow 0$, all energy modes are equally singly occupied and $\lambda=1$. For the isotropic nearest neighbor spin ice, as $\beta\rightarrow\infty$, the lowest two energy modes become flat and fully degenerate and $\lambda\rightarrow\frac{1}{2}$. This tells us that the high-energy modes depopulate, and the degenerate low-energy modes are doubly occupied. For the planar ice, as $\beta\rightarrow\infty$, only one lowest energy mode becomes flat and $\lambda\rightarrow\frac{1}{4}$. If a normal mode becomes critical at any temperature, then $\lambda^{-1}\rightarrow\infty$ or $\lambda\rightarrow 0$. This indicates a phase transition with long-range order in the SCGA.
Figure 7: Lagrange-multiplier for the SCGA. Lagrange-multiplier that fixed the spin constraint in the self-consistent Gaussian approximation. For isotropic spin ice, $\lambda\rightarrow\frac{1}{2}$ as $T\rightarrow 0$. For planar ice, $\lambda\rightarrow\frac{1}{4}$ as $T\rightarrow 0$. For the line-order phase, $\lambda\rightarrow 0$ as $T\rightarrow 0$, suggesting long-range order at $T=0$. For the three remaining phases, the sudden drop of $\lambda\rightarrow 0$ at a finite temperature indicates a finite-temperature phase transition.
The Lagrange multiplier for the six phases is shown in Fig. 7. We can observe the expected behaviour as $T\rightarrow 0$ for isotropic spin ice ($\lambda\rightarrow\frac{1}{2}$) and for planar ice ($\lambda\rightarrow\frac{1}{4}$). For the line-order phase, we can observe how $\lambda$ monotonically decreases as $T$ approaches 0. This suggests a transition to long-range order to appear only at $T=0$, similarly to what is expected in a one-dimensional Ising chain [88]. By contrast, the Lagrange multiplier of omni-plane phase, the plane-paramagnetic phase, and the ferromagnetic phase seem to have a sudden drop to zero at a finite temperature. It is important to note, however, that $\lambda$ can only really reach zero at a finite temperature for $L\rightarrow\infty$. While for the plane-paramegnetic phase and the ferromagnetic phase, the drop in $\lambda$ appears sudden and steep, $\lambda$ of the omni-planar phase approaches small values more gradually. A finite-size scaling analysis of $\lambda$ in the omni-planar phase, however, still points towards a finite-temperature phase transition within the SCGA. This supports the numerical findings in Sec. IV.2, indicating a finite-temperature phase transition in the three phases with any $\delta_{i}<0$.
@@
Appendix E Monte-Carlo simulation
@@
During each Monte-Carlo time step, six types of updates were employed to enhance equilibration: single spin flips, worm loop updates, and other cluster-updates (see App. F). For each measurement temperature, the system was evolved until $N_{\text{sweeps}}=N_{\text{spin}}\cdot 10^{3}$ accepted updates (including spin flips and worm loops) were performed. Measurements were taken at equilibrium.
To generate data across different temperatures, the system was progressively cooled in 10 steps from the last measurement temperature. At each temperature step, the system was equilibrated before proceeding to the next step. Finally, at the next measurement-temperature, an additional equilibration phase was conducted.
Here again, the next sweep measurement was performed.
Appendix F Cluster update schemes
For isotropic nearest-neighbor spin ice, in numerical simulations, typically updates are given by single spin flips as well as worm loop updates [55, 56, 57]. Since at low temperatures, spin freezing makes equilibration times with only single-spin flips extremely long, the worm loop algorithm finds update moves at a zero energy cost [14]. The ground state degeneracy of all anisotropic systems is, however, greatly reduced compared to the isotropic spin ice case. This also leads to an increase in equilibration time in numerical simulations when only employing the two aforementioned update schemes. The problem lies in the acceptance rate of the attempted updates. A measurement series at a particular temperature is only completed if a prefixed number of spin updates are accepted. For isotropic spin ice below the crossover, where the system is in a ground state configuration, all attempted worm loop updates are accepted. However, this acceptance rate drops significantly for all anisotropic systems. Now, to achieve the same number of accepted updates, either the number of attempted updates must be increased significantly, or a new type of update with a higher acceptance rate must be proposed. The latter is oped for to produce the data in Fig. 3 and the particular update schemes are discussed here.
@@ Appendix F Cluster update schemes
The underlying ordering processes are not changed by the introduction of the new cluster-update schemes, but rather, achieving the result numerically is merely accelerated. This is shown exemplarily in Fig. 9 for a system size of $L_{x}\times L_{y}\times L_{z}=8$, where the data obtained by only using the worm loop updates and single spin flips, and all six cluster update schemes detailed above are shown. The two curves overlap fully.